Nadia Corporation adjusts its debt so that its interest coverage (EBIT/Interest) remains constant at 3. Nadia’s EBIT next period is projected to be $15 million and this is expected to grow at 3.5% annually. Nadia expects its incremental capital expenditure in future (net of depreciation) to equal 1% of EBIT every year. Its investment in incremental net working capital is expected to be 1% of EBIT every year and it faces 40% tax rate.

  1. If the unlevered cost of capital for Nadia is 8%, what is its unlevered value?
  2. What is the levered value of Nadia?
  3. If Nadia’s cost of debt is 6% how much debt should it have to begin with?
Answer #1

(a) Expected EBIT = $ 15 million, Tax Rate = 40 % , Net Cap Expenditure = 1% of EBIT and Net Working Capital (NWC) = 1% of EBIT, Perpetual EBIT Growth Rate = 3.5 %

As both Cap Expenditure and NWC are 1% (constant %) of the EBIT, they too shall grow at 3.5 % in perpetuity

Expected FCFF = EBIT x (1-Tax Rate) - Net Cap Expenditure - NWC = 15 x (1-0.4) - 0.01 x 15 - 0.01 x 15 = $ 8.7 million

Unlevered Cost of Capital = 8%

Unlevered Firm Value = 8.7 / (0.08 - 0.035) = $ 193.33 million

(b) EBIT / Interest Expense = 3 and EBIT = $ 15 million

Interest Expense = 15 / 3 = $ 5 million and Tax Rate = 40 %  

Interest Tax Shield = Interest Expense x Tax Rate = 5 x 0.4 = $ 2 million

Cost of Debt = 6 %

Present Value of Interest Tax Shield = 2 / 0.06 = $ 33.33 million

Levered Firm Value = 193.33 + 33.33 = $ 226.66 million

(c) Cost of Debt = 6 % and Interest Expense = $ 5 million

Debt = Interest Expense / Cost of Debt = 5 / 0.06 = $ 83.33 million